# SADG: Segment Any Dynamic Gaussian Without Object Trackers

### v1

@@ 4 Experimental Results
We evaluate the performance of SADG across several well-known datasets to validate its effectiveness. We begin by introducing the datasets in Sec. 4.1. Since there are no established benchmarks for validating the segmentation performance on novel views in dynamic scenes, we provide details on how we extend these different datasets with segmentation benchmarks (Sec. 4.2). Next, we explore the implementation details and various configurations used to train and test methods on these benchmarks (Sec. 4.3). We present segmentation results of our model, along with both quantitative and qualitative comparisons to other related works [6, 57, 25] in Sec. 4.5. Last but not least, we demonstrate capabilities of our learned semantic representation on a number of downstream tasks including scene editing and selection in our a graphical user interface (Sec. 4.6).
4.1 Datasets
@@
SADG (Ours) | 0.8719 | 0.9826 | 0.8663 | 0.9845 | 0.9022 | 0.9945 | 0.9234 | 0.9945 | 0.9308 | 0.9917 |
NeRF-DS dataset [52] contains scenes with challenging transparent and shiny moving objects recorded by two monocular cameras. Five sequences are used for reconstruction and evaluation including as_novel_view, basin_novel_view, cup_novel_view, press_novel_view, and plate_novel_view.
HyperNeRF dataset [36] is recorded using a single monocular camera . The camera moves while capturing the dynamic scenes involving different human-object interactions. In our experiments, we use ten sequences: americano, split-cookie, oven-mitts, espresso, chickchicken, hand1-dense-v2, torchocolate, slice-banana, keyboard, and cut-lemon1.
Neu3D dataset [27] contains videos captured by a multi-view rig with to cameras each at 30 FPS. The scenes involve a person performing cooking tasks in the kitchen. In total, five scenes are used in our experiments: coffee_martini, cook_spinach, cut_roasted_beef, flame_steak, and sear_steak.
Google Immersive dataset [2] consists of videos captured by a 46-camera rig at 30 FPS. The cameras are fish-eye cameras and are distributed on the surface of a hemispherical, 92cm diameter dome. Four scenes are used in our experiments: 01_Welder, 02_Flames, 10_Alexa_1, and 11_Alexa_2.
Technicolor Light Field dataset [44] captures scenes using a camera array at 30 FPS. We use the undistorted images provided by the authors, and four scenes are used in our experiments: Birthday, Fabien, Painter, and Theater.
4.2 Segmentation Benchmarks
To date, while numerous benchmarks exist for dynamic video segmentation, such as DAVIS [38, 40, 3, 4] and VOS [51, 54, 55], there are still no well-established segmentation benchmarks for dynamic novel view synthesis.
Such benchmarks are crucial for evaluating a comprehensive understanding of scenes, both geometrically and semantically. Existing datasets, such as Replica [47], ScanNet [10], and ScanNet++ [58], include ground-truth labels for evaluating panoptic segmentation performance, but they do not cover dynamic sequences. To address this gap, we manually annotate dynamic sequences in our experiments and propose the NeRF-DS-Mask, HyperNeRF-Mask, Neu3D-Mask, Immersive-Mask, and Technicolor-Mask benchmarks. Utilizing the powerful SAM2 model [43], which provides consistent object masks across video sequences, we manually define the objects of interest for each sequence to evaluate their segmentation performance in our model. Examples of these mask objects can be seen in the supplementary.
4.3 Implementation Details
We employ SAM [20] to generate the anything-masks for Gaussian feature learning. However, the resulting masks are memory-inefficient because they are stored in a dense format () with a lot of pixels with zero values. To optimize memory usage, we save the masks as a bit array.
Given a permuted tensor version of an RGB image , SAM generates masks . We first flatten , then convert flattened ()-dimensional array into a bit-array. The final output is stored in a dictionary with the required information to re-create the original mask tensor: a number of masks , image dimensions , and the compressed bit array, i.e., . This approach significantly reduces the storage used and is particularly useful for multi-view datasets, where numerous masks from different views are generated. For instance, a raw mask generated from an image in coffee_martini occupies around 80 MB, but the dictionary representation is significantly reduced to approximately 1 MB. Due to space limits, we provide training details for different datasets in the supplementary.
@@ 4.4 Evaluation Protocol and Baselines
Given ground truth masks we use Mean Intersection over Union (mIoU) and Mean Pixel Accuracy (mAcc) to quantitatively evaluate segmentation performance of the models. We compare our model with SAGA [6], Gaussian Grouping [57] and DGD [25]. Note that both SAGA and Gaussian Grouping are designed for static scenes and do not include a module to encode temporal information. To address this limitation, we integrate our 4D reconstruction backbone into their pipeline, extending their approach to effectively handle dynamic scenes. Moreover, Gaussian Grouping needs the precomputed object IDs from DEVA [8], whose labels are inconsistent for multi-view camera setup. Thus, we train their semantic module on a monocular stream from the camera closest to the test view. Furthermore, all reported baselines utilize the same pre-trained 4D reconstruction, with semantics learned according to each pipeline. This allows us to isolate and evaluate the performance of each model based on its semantic learning capabilities. After training, we manually perform click prompts in the rendered novel views of each scene, which are trained by different approaches to select the objects of interest defined in the benchmark. Apart from the quantitative results, we also present qualitative results from various benchmarks, as certain insights cannot be fully captured through numerical evaluations alone. Since the official source code for SA4D [16] is not available, we report the comparison based on unofficial implementation in the supplementary.
@@ 4.5 Segmentation Results
As shown in Tab. 1, SADG consistently outperforms all other methods across all benchmarks in terms of average mIoU and mAcc. We also select some sequences to visually illustrate the quality of the segmentation of each model in Fig. 5. These comparisons reveal that SADG qualitatively surpasses existing models and exhibits multi-view consistent segmentation performance. Notably, SADG maintains strong generalization to test camera views in multi-view sequences such as flame_steak, Painter, and 11_Alexa_2. This robust performance can be attributed to the compact Gaussian features learned through the proposed Gaussian feature training. In contrast, DGD segments objects by comparing cosine similarity scores between features retrieved via click prompts. However, determining a suitable threshold that generalizes across various scenes is neither straightforward nor intuitive.
SADG employs clustering based on the Gaussian features to group the Gaussians into several segments. As this operation is done in 3D space, SADG achieves cross-view consistency and does not need to associate objects from different views. This advantage is evident in cut-lemon1 example in Fig. 5, where wrongly associated object IDs can cause segmentation failure for Gaussian Grouping.
4.6 Scene Editing Applications
@@
In addition to accurate segmentation capabilities and multi-view consistent rendered masks, we demonstrate the effectiveness of our learned semantic feature field in a range of downstream tasks. Specifically, we focus on interactive editing applications and develop a graphical user interface (GUI) to grant full user control over the representation. Thanks its explicit nature and our semantic abstraction to Gaussian clusters, we achieve real-time operation capability in handling many editing tasks. We provide snapshots of the GUI in the supplementary.
@@
Besides simple click prompts in the novel view of dynamic scenes for object selection, we also support text-based prompts to make object selection more intuitive. To enhance 3D content editing in dynamic scenes, we integrate style transfer capabilities for the selected objects. By segmenting objects directly in 3D, we facilitate scene composition and object removal through straightforward manipulation of the selected Gaussians. Due to space limit, we shall provide more qualitative results in the supplementary. In the following we describe the implementation details of editing tasks based on our semantically-aware feature field.
@@
Text Prompt. Inspired by Gaussian Grouping [57], we employ Grounding DINO [28] to generate 2D mask based on the text prompt in the novel view. The 2D mask is reprojected into the 3D space with the rendered depth information. Finally, we associate each reprojected point with its nearest Gaussian in the 3D scene and retrieve its corresponding cluster. If the number of associated Gaussian of a given cluster is more than a threshold, the cluster would be selected. The example of text prompts on different sequences is shown in Fig. 6 (a).
@@
Scene Composition. We also perform qualitative results on scene composition. As the segmentation is performed in 3D, we can simply put the selected Gaussians from the dynamic scene into another static or dynamic scene. The scales and coordinate transformation need to be adapted. We show some examples on scene composition in Fig. 6.
@@
Object Removal. We can also remove object by simply deleting Gaussians belonging to the specific cluster as illustrated in Fig. 6 (d). It is worth noting that no inpainting is required due to the underlying 3D representation, where exposed parts are learned by multi-view observations.
@@ 5 Conclusion
We introduced a novel framework for dynamic scene understanding, which enables multi-view consistent segmentation without any object tracking supervision. Our SADG effectively combines dynamic 3D Gaussian Splatting 3DGS [56] and 2D SAM [20] masks in a contrastive learning objective, which lifts semantic information into 3D space and learns expressive Gaussian features based on hard pixel cases. This results in cross-view consistency when rendering segmented objects, enhancing the quality and coherence of object segmentation across different views. Evaluated on various novel-view datasets, SADG shows superior performance both quantitatively and qualitatively. We further demonstrated the effectiveness of the learned feature field on downstream editing tasks such as point and text prompts, style transfer, object removal, and scene composition. Our approach sets a strong foundation for further research into dynamic scene understanding and scene editing, especially in complex multiview scenarios.
@@
In the supplementary, we provide further implementation details in Sec. 6 and Sec. 7. A range of ablation studies on different components of our pipeline is given in Sec. 8. We compare our SADG with the concurrent work on 4D segmentation and editing, Segment Any 4D Gaussians (SA4D), in Sec. 9. Several examples from our segmentation benchmark are given in Sec. 11. We present detailed quantitative results on all benchmarks in Sec. 10. We finally provide qualitative results of our method on segmentation (Sec. 12, Sec. 14) and editing (Sec. 13).
@@ 6 Training Implementation Details
We implement the whole pipeline in PyTorch [37]. We reuse the codebase from Deformable-3DGS [56] and extend the differential Gaussian rasterization pipeline to render the Gaussian features similar to [57]. In the following, we introduce the detailed training setup for different datasets.
@@ 6.1 NeRF-DS and HyperNeRF
SADG (w/o filtering) | 0.8719 | 0.9826 | 0.8368 | 0.9790 | 0.8738 | 0.9923 | 0.9154 | 0.9942 | 0.9258 | 0.9904 |
Ours (w/ filtering) | 0.8719 | 0.9826 | 0.8663 | 0.9845 | 0.9022 | 0.9945 | 0.9234 | 0.9945 | 0.9308 | 0.9917 |
Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
SADG (K=0) | 0.8228 | 0.9706 | 0.8360 | 0.9787 | 0.8661 | 0.9920 | 0.8725 | 0.9909 | 0.9259 | 0.9908 |
Ours (K=16) | 0.8719 | 0.9826 | 0.8368 | 0.9790 | 0.8738 | 0.9923 | 0.9154 | 0.9942 | 0.9258 | 0.9904 |
SADG (K=32) | 0.8754 | 0.9828 | 0.8331 | 0.9780 | 0.8702 | 0.9922 | 0.9217 | 0.9944 | 0.9255 | 0.9904 |
@@
6.2 Neu3D

### v2

@@ 4 Experimental Results
We first provide details of our proposed 4D segmentation benchmark in Sec. 4.1 and outline the evaluation protocol in Sec. 4.2 . We present segmentation results of our model, along with both quantitative and qualitative comparisons to other related works [6, 53, 24] in Sec. 4.3. Last but not least, we demonstrate the capabilities of our learned semantic representation on a number of downstream tasks, including scene editing and selection via our designed graphical user interface (Sec. 6). Please refer to the supplementary material for implementation details.
4.1 4D Segmentation Benchmark
We manually annotate dynamic sequences of existing novel-view synthesis datasets and present their segmentation-aware versions: NeRF-DS-Mask, HyperNeRF-Mask, Neu3D-Mask, Immersive-Mask, and Technicolor-Mask. Utilizing the powerful SAM2 model [38], which provides consistent object masks across video sequences, we define the objects of interest for each sequence to evaluate their segmentation performance in our model and manually refine them. Qualitative examples of the augmented sequences, together with a description of the raw datasets, can be found in the supplementary material.
@@
| SA4D [19] | 0.6740 | 0.9360 | 0.7371 | 0.9616 | 0.8832* | 0.9391* | 0.7987* | 0.9835* | 0.8271* | 0.9734* |
| CGC [42] | 0.8100 | 0.9718 | 0.8157 | 0.9768 | 0.8754 | 0.9927 | 0.8856 | 0.9925 | 0.8806 | 0.9812 |
| OpenGaussian [46] | 0.6939 | 0.9564 | 0.6311 | 0.9411 | 0.8178 | 0.9899 | 0.7120 | 0.9793 | 0.8390 | 0.9694 |
| TRASE (Ours) | 0.8768 | 0.9831 | 0.8663 | 0.9845 | 0.9022 | 0.9945 | 0.9234 | 0.9945 | 0.9308 | 0.9917 |
@@ 4.2 Evaluation Protocol and Baselines
We compare our model with SAGA [6], Gaussian Grouping (GG) [53], SA4D [19], and DGD [24]. Additionally, we add evaluation with the contrastive learning methods such as Contrastive Gaussian Clustering (CGC) [12] and OpenGaussian [46]. Note that SAGA, Gaussian Grouping, CGC, and OpenGaussian are designed for static scenes and do not include a module to encode temporal information. To evaluate them on the 4D segmentation task, we integrate our 4D reconstruction backbone into their pipeline, extending their approach to effectively handle dynamic scenes. GG and SA4D require video object tracking labels for their supervision. To enable evaluation of the multi-view sequences from our benchmarks, we train the models on the camera closest to the test view (cam5 for Neu3D, cam2 for Immersive, and cam1 for Technicolor). After training, we manually perform click prompts in the rendered novel views of each scene trained by different approaches to select the objects of interest defined in the benchmark. Apart from the quantitative results, we also present qualitative results from various benchmarks, as certain insights cannot be fully captured through numerical evaluations alone. Following common practices, we use Mean Intersection over Union (mIoU) and Mean Pixel Accuracy (mAcc) to quantitatively evaluate the performance of the models.
@@ 4.3 Segmentation Results
| Preproc | Training | Total (min) | Storage (MB) | |
| DGD [24] | 0.08 | 103 | 103 | 285 |
| TRASE (Ours) | 16 | 16 | 32 | 60 |
@@
As shown in Tab. 1, TRASE achieves state-of-the-art performance across all benchmarks, surpassing other segmentation methods. Please refer to the supplementary material for details on each sequence of the benchmark. Further, we select some sequences to illustrate the quality of the segmentation of each model in Fig. 5. These comparisons reveal that TRASE qualitatively surpasses existing models and exhibits multi-view consistent segmentation performance. Notably, TRASE maintains strong segmentation consistency to test camera views in multi-view sequences such as flame_steak, Painter, and 11_Alexa_2. This robust performance can be attributed to the compact Gaussian features learned through the proposed Gaussian feature training. In contrast, DGD segments objects by comparing cosine similarity scores between features retrieved via click prompts. However, determining a suitable threshold that generalizes across various scenes is neither straightforward nor intuitive. TRASE employs clustering based on the Gaussian features to group the Gaussians into several segments. As this operation is done in 3D space, TRASE achieves cross-view consistency and doesn’t need to associate objects from different views. This advantage is evident in cut-lemon1 example in Fig. 5, where wrongly associated object IDs can cause segmentation failure for [53].
5 Ablation Studies
5.1 Time and Storage Analysis
We perform a comprehensive time and storage analysis of TRASE and DGD [24] using an RTX 3080 and an Intel i7-12700 on the NeRF-DS [48] dataset. The results are shown in Tab. 2, where “Pre-Proc” refers to the generation of DINOv2 384-dim features for DGD and SAM masks for TRASE. Rendering 384-dim features into 2D is computationally inefficient, significantly slowing down DGD’s training process. TRASE uses 32-dim Gaussian features, which can be rendered much more efficiently. As a result, TRASE achieves over 3 faster total time than DGD while only requiring 4.5 less space in storage.
5.2 Soft Sample Mining
We visually demonstrate the effectiveness of our soft-mined pixel-pair sampling strategy in comparison to other contrastive objectives: sampling all positive–negative pairs (all) and sampling only pairs that individually satisfy the thresholds (hard). We include contrastive-based baselines such as CGC [12] and OpenGaussian [46] for completeness. By investigating the out-of-FoV regions, i.e., 3D regions outside the observable space of train and test views, we observed floaters in the segmentation by our method without soft selection and other contrastive baselines, as shown in Fig. 6. Our mining approach effectively eliminates these floaters, leading to more stable and reliable segmentation. It is worth noting that no post-processing filtering is applied to ensure a fair comparison.
5.3 Single-frame Mask Supervision
Inspired by classical tracking methods [15] and recent work linking 3DGS dynamics to pixel motion [14], we investigate the impact of incorporating temporal constraints on our single-frame mask supervision. Specifically, we compare our standard single-frame approach with a temporal extension, where masks are propagated across consecutive frames to encourage segmentation consistency over time. To implement this, we estimate optical flow between consecutive training timestamps and using differentiable rendering with accumulated 2D Gaussian displacements [14], and warp mask pixels from to . Quantitative results are presented in Tab. 3. As observed, the multi-frame baseline does not provide significant improvements over the single-frame supervision, despite the doubled training cost, as an additional 2D flow and all associated parameters are rendered. This validates our design choice to rely on single-frame mask supervision, which achieves strong segmentation performance efficiently.
| HyperNeRF-Mask | Immersive-Mask | Technicolor-Mask | ||||
| Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
| TRASE (multi-frame) | 0.8636 | 0.9833 | 0.8968 | 0.9933 | 0.9309 | 0.9913 |
| TRASE (Ours) | 0.8663 | 0.9845 | 0.9234 | 0.9945 | 0.9308 | 0.9917 |
| NeRF-DS-Mask | HyperNeRF-Mask | Neu3D-Mask | |||||
| Method | Filtering | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
| SAGA [6] | ✘ | 0.8072 | 0.9644 | 0.7660 | 0.9579 | 0.6941 | 0.9516 |
| ✓ | 0.8421 | 0.9778 | 0.8223 | 0.9779 | 0.8026 | 0.9853 | |
| TRASE | ✘ | 0.8760 | 0.9829 | 0.8368 | 0.9790 | 0.8738 | 0.9923 |
| TRASE (Ours) | ✓ | 0.8768 | 0.9831 | 0.8663 | 0.9845 | 0.9022 | 0.9945 |
5.4 Post-processing Filtering
We provide an ablation study of our method with respect to the filtering step, which removes Gaussians whose features are too dissimilar to the object’s mean. As shown in Tab. 4, the proposed filtering step enhances the segmentation quality of our method. Particularly, artifacts on the object’s boundaries are effectively removed, mitigating the inaccuracies in input masks or the effects of occlusions.
We further demonstrate the advantage of our learned feature field by showing that while the post-processing step enhances segmentation accuracy, it is not the sole factor driving our superior performance. We also note that this technique is applicable only to certain baselines such as SAGA [6]. For example, methods like CGC [12] and DGD [24] already incorporate a similarity threshold during inference. OpenGaussian [46] gets the assigned class ID for each Gaussian from their codebook. By contrast, SA4D [19] and Gaussian Grouping [53] learn an identity field, making our post-processing inapplicable to them. In Tab. 4 and Fig. 7, we quantify and illustrate the effectiveness of applying our proposed filtering step to SAGA [6]. However, because the features learned under the baseline’s hard-contrastive objective remain suboptimal, the segmentation quality still exhibits more artifacts compared to TRASE.
6 Editing Applications
@@
In addition to accurate segmentation capabilities and multi-view consistent rendered masks, we demonstrate the effectiveness of our learned semantic feature field in a range of downstream tasks. Specifically, we focus on interactive editing applications and develop a graphical user interface (GUI) to grant full user control over the representation. Thanks to its explicit nature and our semantic abstraction to Gaussian clusters, we achieve real-time operation capability in handling many editing tasks. We provide snapshots of the GUI in the supplementary.
@@
In addition to simple click prompts, we also support text-based prompts to make object selection even more intuitive. To enhance 3D content editing in dynamic scenes, we integrate style transfer capabilities for the selected objects. By segmenting objects directly in 3D, we facilitate scene composition and object removal through straightforward manipulation of the selected Gaussians. Due to the space limit, we shall provide more qualitative results in the supplementary. In the following, we describe the implementation details of editing tasks based on our semantically-aware feature field.
@@
Text Prompt. Inspired by Gaussian Grouping [53], we employ Grounded SAM [39] to generate a 2D mask based on the text prompt in the novel view. The 2D mask is reprojected into the 3D space with the rendered depth information. Finally, we associate each reprojected point with its nearest Gaussian in the 3D scene and retrieve its corresponding cluster. If the number of associated Gaussians of a given cluster is more than a threshold, the cluster would be selected. The example of text prompts on different sequences is shown in Fig. 8 (a).
@@
Scene Composition. We also perform qualitative results on scene composition. As the segmentation is performed in 3D, we can simply put the selected Gaussians from the dynamic scene into another static or dynamic scene. The scales and coordinate transformation need to be adapted. We show some examples of scene composition in Fig. 8.
@@
Object Removal. We can also remove objects by simply deleting Gaussians belonging to the specific cluster as illustrated in Fig. 8 (d). It is worth noting that an additional inpainting step may be required, where the newly exposed parts have not been learned from multi-view observations.
@@ 7 Conclusion
We introduced a novel framework for dynamic scene understanding, which enables multi-view consistent segmentation without any object tracking supervision. Our TRASE effectively combines dynamic 3D Gaussian Splatting 3DGS [52] and 2D SAM [22] masks in a contrastive learning objective, which lifts semantic information into 3D space and learns expressive Gaussian features based on soft sample mining. This results in cross-view consistency when rendering segmented objects, enhancing the quality and coherence of object segmentation across different views. Evaluated on various novel-view datasets, TRASE shows superior performance both quantitatively and qualitatively. We further demonstrate the effectiveness of the learned feature field on downstream editing tasks such as point and text prompts, style transfer, object removal, and scene composition. Our approach sets a strong foundation for further research into dynamic scene understanding and scene editing, especially in complex multiview scenarios. We intend to release our code and benchmarks for future developments.
@@
In the supplementary, we provide several examples from our segmentation benchmark, along with dataset details (Sec. 8). Further implementation details are described in Sec. 9 and Sec. 10. A range of ablation studies on different components of our pipeline is given in Sec. 11. In Sec. 12 we address several limitations of our method and potential directions for future work. We provide additional visualizations of our learned feature space in comparison to other contrastive-based segmentation methods in Sec. 13 and Sec. 14. In Sec. 15 we show more examples on how our soft mining successfully mitigates floaters. Further, we present detailed quantitative results on all benchmarks in Sec. 17. We finally demonstrate qualitative results of our method on segmentation (Sec. 18, Sec. 19) and editing (Sec. 16).
8 Segmentation Benchmark
As discussed in Sec. 4.1 (main), we employ SAM2 [38] to manually select objects of interest to form our segmentation benchmarks for dynamic scene novel view synthesis. The examples are illustrated in Fig. 9. We describe the details of the mask storage for training and datasets adapted for our segmentation benchmark in the following.
Mask Storage. We employ SAM [22] to generate the anything-masks for Gaussian feature learning. However, the resulting masks are memory-inefficient because they are stored in a dense format () with a lot of zero-pixels. To optimize memory usage, we save the masks as a bit array. Given a permuted tensor version of an RGB image , SAM generates masks . We first flatten , then convert flattened ()-dimensional array into a bit-array. The final output is stored in a dictionary with the required information to re-create the original mask tensor: a number of masks , image dimensions , and the compressed bit array, i.e., . This approach significantly reduces the storage used and is particularly useful for multi-view datasets, where numerous masks from different views are generated. For instance, a raw mask generated from an image in coffee_martini occupies around 80 MB, but the dictionary representation reduces to about 1 MB.
NeRF-DS. [48] contains scenes with challenging transparent and shiny moving objects recorded by two monocular cameras. Five sequences are used for reconstruction and evaluation including as_novel_view, basin_novel_view, cup_novel_view, press_novel_view, and plate_novel_view.
HyperNeRF. [33] is recorded using a single monocular camera
. The camera moves while capturing the dynamic scenes involving different human-object interactions. In our experiments, we use ten sequences: americano, split-cookie, oven-mitts, espresso, chickchicken, hand1-dense-v2, torchocolate, slice-banana, keyboard, and cut-lemon1.
Neu3D. [26] contains videos captured by a multi-view rig with to cameras each at 30 FPS. The scenes involve a person performing cooking tasks in the kitchen. In total, five scenes are used in our experiments: coffee_martini, cook_spinach, cut_roasted_beef, flame_steak, and sear_steak.
Google Immersive. [2] consists of videos captured by a 46-camera rig at 30 FPS. The cameras are fish-eye cameras and are distributed on the surface of a hemispherical, 92cm diameter dome. Four scenes are used in our experiments: 01_Welder, 02_Flames, 10_Alexa_1, and 11_Alexa_2.
Technicolor Light Field. [40] captures scenes using a camera array at 30 FPS. We use the undistorted images provided by the authors, and four scenes are used in our experiments: Birthday, Fabien, Painter, and Theater.
@@ 9 Training Implementation Details
We implement the whole pipeline in PyTorch [34]. We reuse the codebase from Deformable-3DGS [52] and extend the differential Gaussian rasterization pipeline to render the Gaussian features similar to [53]. In the following sections, we present the detailed training and evaluation setup for various datasets.
@@ 9.1 NeRF-DS and HyperNeRF
| TRASE (K=0) | 0.8228 | 0.9706 | 0.8360 | 0.9787 | 0.8661 | 0.9920 | 0.8725 | 0.9909 | 0.9259 | 0.9908 |
| Ours (K=16) | 0.8760 | 0.9829 | 0.8368 | 0.9790 | 0.8738 | 0.9923 | 0.9154 | 0.9942 | 0.9258 | 0.9904 |
| TRASE (K=32) | 0.8754 | 0.9828 | 0.8331 | 0.9780 | 0.8702 | 0.9922 | 0.9217 | 0.9944 | 0.9255 | 0.9904 |
@@
) on the objects of interest in the novel view to segment the target objects.9.2 Neu3D